Calculate the quantities for ideal gas given below.
(R=0.082 L atm mol-1 K-1)
i. The absolute temperature of 21 g Nitrogen gas at which gas occupies 164 mL at 6 atm.
ii. The pressure in atm, if 8.25 X 10 -2 mole gas occupies a volume of 255 mL at 115ºC.
iii. The molecular weight of 12.3 g gas if 5.49 L at 35ºC have pressure 11.25 atm.
iv. A closed gas cylinder contains 0.500 mole H2, 0.300 mole O2, 1.200 mole CO2 at 25 °C and pressures 2.00 atm. Calculate the volume of the cylinder.
v. Calculate the partial pressure of O2 inside the cylinder in question (iv).
1
Expert's answer
2020-07-29T06:38:52-0400

From the equation: PV = nRT, P - pressure (atm.), V - volume (L) R = 0.082 L atm mol-1 K-1, T - temperature (K, 0°C = 273K), n - amount of substance (mol)

1) T = PV/RT = 16 K

2) P = nRT/V = 10,3 atm

3) n = m/M, M = mRT/PV = 5 g/mol

4) n = 0,3 + 0,5 + 1,3 = 2 mol, V = nRT/P = 24,436 L

5) P(O2) = P*n(O2)/n = 2*0,3/2 = 0,3 atm